OpenStudy (anonymous):

A 4.4 kg object is being pushed along a surface, causing it to accelerate at a rate of 1.5 m/s2 . Th e coeffi cient of kinetic friction is 0.25. What is the magnitude of the horizontal force being applied to push the object?

OpenStudy (anonymous):

Sum of the forces = mass * acceleration there for Fnet = 4.4*1.5 = 6.6 N The two forces that are acting on the object are friction (Ff) and the applied force (Fa) Ff = mu*Normal force = mu*m*g = 0.25*4.4*9.8 = 10.78 N Also Fnet = Fa - Ff So Fa = Fnet + Ff = 10.78 +6.6 = 17.38 N I hope that helps

OpenStudy (anonymous):

therefore*

OpenStudy (anonymous):

Thank you so much @melaghil !